The average number of calories in a 1.8-ounce chocolate bar is 230. Suppose


that the distribution of calories is approximately normal with a standard


deviation of 10. Find the probability that a randomly selected chocolate bar will


have less than 200 calories.



1
Expert's answer
2022-03-30T09:21:25-0400

"P(X<200)=P(Z<\frac{200-230}{10})"

"=P(Z<-3)"

From the normal tables

"=0.0013"